OpenStudy (anonymous):

Combine. Assume that all variables represent nonnegative real numbers. square root 75a^3 + a square root 12a

OpenStudy (anonymous):

\[\sqrt{75x^3}=\sqrt{25\times 3\times x^2\times x}=\sqrt{25}\sqrt{x^2}\sqrt{3x}=5x\sqrt{3x}\] is a start